OpenStudy (anonymous):

use the product to sum formula to write the product as a sum or difference 10 cos 75° cos 15°

OpenStudy (raden):

use the identity : 2 cosAcosB = cos(A+B) + cos(A-B)

OpenStudy (anonymous):

thanks! @RadEn

OpenStudy (raden):

yw did u get it ?

OpenStudy (anonymous):

I did

OpenStudy (raden):

that's cool :)

OpenStudy (raden):

so, what ur answer --"

OpenStudy (anonymous):

well I think I got it. isn't is Cos(75+15) + Cos(75-15) which is Cos (90) + Cos (60)

OpenStudy (raden):

then, final answer is ...

OpenStudy (anonymous):

Cos (90) =0 and Cos (60) = 1/2 so 0 +1/2 = 1/2

OpenStudy (raden):

good, but still incorrect, hihik :P

OpenStudy (anonymous):

What am I doing wrong? :O

OpenStudy (raden):

the final answer should be 5/2 :)

OpenStudy (anonymous):

5cos(75 + 15) + 5 cos(75 - 15) the terms 5sin75sin15 -5sin75sin15 = 0

OpenStudy (anonymous):

hmmmm. i'm still not getting that

OpenStudy (raden):

10 cos 75° cos 15° = 5 (2 cos 75° cos 15°) = 5 (you have already got this above) = 5 (1/2) = 5/2 hehe :)

OpenStudy (anonymous):

hahaha thanks, you genius!

OpenStudy (raden):

you're welcome ;)

OpenStudy (anonymous):

just wondering are you a teacher? lol

OpenStudy (anonymous):

jk. you're a junior. .